OpenStudy (anonymous):

How long is a wire reaching from the top of a 12-foot pole to a point 6feet from the base of the pole? Answer

OpenStudy (anonymous):

its a right triangle. draw it

OpenStudy (anonymous):

6^5 6^3 3^5 6

OpenStudy (anonymous):

why is six bigger than twelve?

OpenStudy (anonymous):

oh hush :) i drew it in 10 seconds

OpenStudy (anonymous):

\[\sqrt{12^2+6^2}\] \[\sqrt{144+36}\] \[\sqrt{180}\]

OpenStudy (anonymous):

\[=6\sqrt{5}\] if you want simplest radical form